how do I set up and solve this problem
we purchased 60 units and gave away three and sold the rest at twice their purchase price. The total profit eas $1188 how much did we sell them for?

Answer by ankor@dixie-net.com(22740)

we purchased 60 units and gave away three and sold the rest at twice their purchase price. The total profit was $1188 how much did we sell them for?
;
Let x = purchase price
then
2x = selling price
:
Revenue - cost = profit
2x(60-3) - 60x = 1188
:
2x(57) - 60x = 1188
:
114x - 60x = 1188
54x = 1188
x = 1188/54
x = $22 the purchase price
then
$44 price they sold them for
:
:
Check this:
44(57) - 22(60) =
2508 - 1320 = 1188